### v1

# DMUS-MP-25/04

Abstract
The massless $S$-matrix of the pure RR $AdS_{3}\times S^{3}\times T^{4}$ theory is very similar but not quite the same as Fendley-Intriligator’s ${\cal{N}}=2$ $S$-matrix, in turns related to Sine-Gordon taken at a special coupling. In this short note we review the reason why supersymmetry emerges but with a fractional statistics of the particles. We then use this to obtain a very simple interpolating $S$-matrix between massless $AdS_{3}$ and Sine-Gordon, and further find two interpolating $S$-matrices between these and the mixed-flux relativistic $S$-matrix, solving the Yang-Baxter equation all the way in-between. They are all relativistic invariant and braiding unitary, can be written in terms of particles with fractional statistics, and they might serve the purpose to embed into a parent integrable system.
1 Introduction and Conclusions
It is undeniable that integrability of strings propagating on $AdS_{3}\times S^{3}\times T^{4}$ [1] (for reviews see [2, 3, 4]) continues to offer opportunities for surprises, keeping on with the tradition inaugurated by $AdS_{5}$ and $AdS_{4}$ [5]. The exact $S$-matrix theory is one of the richest available [6, 7, 8, 9, 10, 11] and the presence of massless modes [12, 13, 14], see also [15] and [16, 17], keeps pushing the technology further and further. The Quantum Spectral Curve (QSC) [18], see also [19], is undergoing an expansion of its methods to accommodate this model [20]. Meanwhile [21], see also [22, 23, 24], have considerably revitalised the field with new algebraic data and have created an exciting race with significant progress on all sides [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
A very important feature of this model is the possibility of a deformation in the form of mixed-flux extension [36, 37, 38, 39, 40] which shows up in the dispersion relation of the elementary scattering particles [41]:
$E=\sqrt{\Big{(}m+\frac{k}{2\pi}p\Big{)}^{2}+4h^{2}\sin^{2}\frac{p}{2}}.$ (1.1)
Notably $k\in\mathbbmss${Z} stands for the WZW level and $m$ for is the mass of the particle. An important wealth of work on this and similar deformations has now been done [42]. As it is customary, one calls pure RR the $k=0$ case (and mixed-flux otherwise).
There is a special subarea inside this field which is evolving into its own little garden, and that is the relativistic sector. This goes from the BMN limit being non-trivial for same-chirality massless scattering [14], where it describes a CFT á lá Zamolodchikov [43], to the entire massless sector being of difference form [34, 35, 13, 21], to the mixed-flux scattering being amenable to a variety of massive relativistic limits [35, 23, 24], all pointing toward famous integrable $S$-matrices of the past - but with differences. The observation [14] is that the massless $S$-matrix of the pure RR theory is very similar but not quite the same as Fendley-Intriligator ${\cal{N}}=2$ $S$-matrix [44, 45, 46], in turns related to Sine-Gordon taken at a special coupling. Considerations on these connections in the case of mixed-flux are also expanded upon in [23], and in particular in [24], where the emphasis is on the differences in the global structure of the supersymmetry algebra.
The focus of this paper is instead the difference in the nature itself of the supersymmetry, in particular the appearance of fractional statistics. We review the detailed reason why supersymmetry accidental pops out but is not exactly that which one has in pure RR $AdS_{3}$, due to a different assignment of fermionic number of the particles. We use the coproduct of the underlying quantum affine symmetry of Sine-Gordon to show this, following the discoverer of this phenomenon Bernard and LeClair [47], see also section 2 of [44] and [48]. We use this to obtain a very simple interpolating $S$-matrix between massless $AdS_{3}$ and Sine-Gordon, which could be useful.
Then we felt inspired and deformed this interpolating $S$-matrix to include something like a non special coupling, basically relaxing the special coupling and letting it be generic. The hope was to obtain an interpolation to the (or one of the) mixed flux $S$-matrix (-ces). It did not exactly work like that, but we managed to find a couple of interpolations which do the job anyway, and solve the Yang-Baxter equation (YBE) all the way in-between. They are all relativistic invariant, and satisfy phyical unitarity for a choice of parameters. This means that we have found a few (perhaps new) relativistic integrable systems ($S$-matrices solving YBE) interpolating Fendley-Intriligator, massless $AdS_{3}$ and mixed flux relativistic. They are all regular, namely $S(\theta)\to\mathbbmss{1}\otimes\mathbbmss{1}$ as $\theta\to 0$, which means that they should sit within the classification of [49], possibly section 4.3.1 modulo twists [^1] They are also all of free-fermion type, which is consistent with having all the AdS/CFT $S$-matrix of free-fermion type - and also the Sine-Gordon $S$-matrix at special coupling $\xi\to 2\pi$, since Sine-Gordon satisfies $\mbox{const.}=4\cos^{2}\frac{\pi^{2}}{\xi}$ in formula (2.2) of [30]. They all have a semblance of $\mathfrak{sl}_{2}$-type quantum symmetry, they are braiding unitarity in deference to their quantum group origin, and they can be rephrased in terms of particles with fractional statistics [^2] Statistics in $1+1$ dimension is a fascinating subject, see for instance [50, 33].. Inspired by [33], we highlight this fact by calling them all fractional $S$-matrices - what seems to be a ubiquitous feature at this point [45].
Ultimately it is quite expected that a minimum continuation/deformation of sorts would fill up the gaps between these models. We believe however that it serves a useful purpose to explicitly embed all the relativistic $S$-matrices appearing from the study of string theory in $AdS_{3}$ inside one single parent relativistic model, which can actually also exactly include Fendley-Intriligator, a.k.a. Sine-Gordon at a special value, showing the common root in terms of fractional statistics. We cannot but entertain the hope that our interpolations are perhaps related to the ones which are being pursued in the literature from the worldsheet viewpoint, such as [4, 31, 32, 29] and the very resembling [17], while we are here focused on trying to follow the tracks left by the simulacrum of the Sine-Gordon $S$-matrix. This paper only treats one block (left-left) for lightness, as if we are displaying a principle and not a complete armamentary, but for completeness one should extend to all chirality blocks - deriving an interpolating crossing equation and parity properties will be particularly sensitive to that assemblage. Incorporating bound states [23] will also be important.
It would also be very interesting to ascertain whether there is a natural way to identify the fractional statistics of the excitations within the structure of the QSC [18, 19, 20].
2 Sine-Gordon at special coupling and $AdS_{3}$
Let us begin by demonstrating how the Sine-Gordon model taken at a special value of the coupling develops an ${\cal{N}}=2$ supersymmetry. More precisely, the statement which we will prove - and which was first observed by [47] - is the following: at the special coupling $\beta^{3}=\frac{16\pi}{3}$ the usual quantum group symmetry of the Sine-Gordon model becomes isomorphic to an ${\cal{N}}=2$ supersymmetry in the soliton-antisoliton sector [^3] As we will shortly review, this is completely internal to Sine-Gordon itself, spontaneously emergent so to say. This is not what is often called ${\cal{N}}=2$ supersymmetric Sine-Gordon (references can be found for instance in [14]), which is instead an extension of Sine-Gordon including more fields to make the Lagrangian supersymmetric to start with. In our case, the Lagrangian is the good old Sine-Gordon, and the supersymmetry happens as an accident for the special value of $\beta$., with fractional fermionic number assignment of the states $|\phi|=\mbox{soliton}=\frac{1}{2}$ and $|\psi|=\mbox{antisoliton}=-\frac{1}{2}$.
Let us recall that the soliton-antisoliton Sine-Gordon $S$-matrix (for a review see for instance [51]) takes the form
$S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&S_{R}&S_{T}&0\\ 0&S_{T}&S_{R}&0\\ 0&0&0&1\end{pmatrix}$, (2.1)
where
$S_{T}=\frac{\sinh\frac{\pi\theta}{\xi}}{\sinh\frac{\pi(i\pi-\theta)}{\xi}},\qquad S_{R}=\frac{i\sin\frac{\pi^{2}}{\xi}}{\sinh\frac{\pi(i\pi-\theta)}{\xi}},\qquad\theta=\theta_{1}-\theta_{2},$ (2.2)
written with the in-basis being $\{|\phi\rangle\otimes|\phi\rangle,|\phi\rangle\otimes|\psi\rangle,|\psi\rangle\otimes|\phi\rangle,|\psi\rangle\otimes|\psi\rangle$}. The variable $\theta_{i}$ is the relativistic rapidity of the $i$-th particle, and $\xi=\frac{\beta^{2}}{8}\frac{1}{1-\frac{\beta^{2}}{8\pi}}$.
We have made the choice in this article to always call the states
$|\phi\rangle=\begin{pmatrix}1\\ 0\end{pmatrix},\qquad|\psi\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$, (2.3)
irrespective or their fermionic grading in different pictures/theories. The context should hopefully help assigning the correct grading, in fact we will try to explicitly say it every time. This is to unclutter the notation.
Proof of fractional supersymmetry
The soliton-antisoliton Sine-Gordon $S$-matrix has quantum affine symmetry $U_{q}\big{(}\widehat{\mathfrak{sl}}(2)\big{)}$ [47]. To start, let us focus on one particular generator $\hat{E}$, such that
$\Delta_{21}\hat{E}\,\,S=S\,\,\Delta\hat{E}$, (2.4)
with
$\Delta\hat{E}=\hat{E}\otimes\mathbbmss{1}+q^{-2H}\otimes\hat{E}$, (2.5)
and where $\hat{E}$ is represented by the matrix $x_{i}E_{\phi\psi}$, having set $x_{i}=\exp[\frac{\pi\theta_{i}}{\xi}]$, for the particle $i=1,2$. $E_{\rho\tau}$ is the matrix with only one non-zero entry equal to $1$ in row $\rho$, column $\tau$. We also have $q=\exp\Big{[}\frac{8\pi^{2}i}{\beta^{2}}\Big{]}$, and $H=\frac{1}{2}(E_{\phi\phi}-E_{\psi\psi})$.
At the special supersymmetric value $\beta^{2}=\frac{16\pi}{3}$ we see that q=iq=-i. The $S$-matrix reduces to
$S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$. (2.6)
We also have
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(\hat{E}\otimes\mathbbmss{1}+q^{-2H}\otimes\hat{E})|\phi\rangle\otimes|\psi\rangle=0+(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle,$ (2.7)
Both $|\phi\rangle$ and $|\psi\rangle$ have the same fermionic number $|\phi|=|\psi|=0$ in Sine-Gordon, and therefore $\hat{E}$ has zero fermionic number - because it sends the state $|\psi\rangle$ into the state $|\phi\rangle$ [^4] The fact that soliton and anti-soliton Bethe eigenstates behave like fermions, even though the bare statistics of the vectors in this original picture is 0, is due the the Sine-Gordon $S$-matrix satisfying $S_{\rho\rho}(0)=-1$, $\rho=\phi,\psi$ [52]..
We can now rephrase (2.7) in a new (fractional) picture:
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\phi\rangle\otimes|\psi\rangle=(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle+0.$ (2.8)
Expressions (2.7) and (2.8) giving the same numerical vector is achieved by this: first, $\hat{E}$ is represented as $x_{i}E_{\phi\psi}=x_{i}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$, but has fermionic number $|\hat{E}|=1$; second, $|\phi\rangle$ is represented as $\begin{pmatrix}1\\ 0\end{pmatrix}$, but has fermionic number $|\phi|=\frac{1}{2}$; third, $|\psi\rangle$ is represented as $\begin{pmatrix}0\\ 1\end{pmatrix}$, but has fermionic number $|\psi|=-\frac{1}{2}$. This means that
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle=(-)^{|\hat{E}||1|}|\phi\rangle\otimes\hat{E}|\psi\rangle,$ (2.9)
which gives the same result.
This exercise can be repeated for a partner supercharge $\hat{F}=x_{i}\,E_{\psi\phi}$, $\Delta\hat{F}=\hat{F}\otimes\mathbbmss{1}+q^{2H}\otimes\hat{F}$. We shall not need the entire ${\cal{N}}=2$ algebra for our scopes, but just half will suffice. The extension of this argument to all the other generators and all other states follows. This reveals the emergence of supersymmetry at this special value of the coupling: we have managed to rephrase the action of the coproduct of the quantum group generators into an equivalent form involving a supercharge, which consistently acts on states with fermionic number $\pm\frac{1}{2}$ in the (graded) tensor product. The coproduct action has become local - as exemplified by (2.8).
In some sense all that has happened is that the quantum-group deformation $q^{-2H}$, resp. $q^{2H}$, in the coproduct at this value of the coupling become equal to $(-1)^{|\hat{E}|\begin{pmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{pmatrix}}$, resp. $(-1)^{|\hat{F}|\begin{pmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{pmatrix}}$, which simulate the statistical factors which a fractional supersymmetry would produce when swapping the corresponding generators with the states. We have again opted for using the same symbol for the generators represented by the same matrices, despite the fact that they might have different fermionic grading, again hoping that the distinction will always be clear and to lighten the notation. This way it will always be
$\hat{E}\propto E_{\phi\psi},\qquad\hat{F}\propto E_{\psi\phi}.$ (2.10)
2.1 Relation with massless pure-RR $AdS_{3}\times S^{3}\times T^{4}$
The Sine-Gordon $S$-matrix at the special value of the coupling discussed in the previous section reduces to a known $S$-matrix discussed by Fendley and Intriligator [44, 45, 46]. This $S$-matrix is almost the same as the massless [^5] For definiteness, we shall always mean left-left same-chirality massless scattering for the sake of this paper. $AdS_{3}\times S^{3}\times T^{4}$, but not quite. It turns out that it differs from it by some factors of $\pm i$ [14], and this is because it has the exact same ${\cal{N}}=2$ supersymmetry which we have just described, but the fermionic assignement of the states is different: in fact in $AdS_{3}$ the fermionic number is the more traditional $|\phi|=0$, and $|\psi|=-1$ (or, equivalently, $|\psi|=1$ mod 22) [^6] This fermionic number is the one used in [53], which therefore ends up being exactly like massless $AdS_{3}$ although apparently for the wrong reason (see comment at the very end of section 3 of [45]).. Let us summarise here below the situation - notice that we use everywhere in this paper the symbol $\theta_{i}$ for the particle’s relativistic rapidity, even if in the case of massless pure RR $AdS_{3}$ we mean the relativistic $\gamma_{i}$ variable [34]. Without any index, $\theta$ always indicates the rapidity-difference.
Except where it would cause an obvious confusion, we have made the choice to call all the $S$-matrices $S$ and all the dressing phases $\Phi$ regardless of the different cases and models, to avoid a needless proliferation of symbols.
  • Fendley-Intriligator:
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.11)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{F},\qquad|\phi|=\frac{1}{2},\quad|\psi|=-\frac{1}{2},\quad|\hat{E}|=1,\quad|\hat{F}|=-1,$
    with $\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi}$.
  • Right-right massless $AdS_{3}\times S^{3}\times T^{4}$:
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.12)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+\mathbbmss{1}\otimes\hat{F},\quad|\phi|=0,\qquad|\psi|=-1,\quad|\hat{E}|=1,\quad|\hat{F}|=-1,$
    and again $\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi}$.
The difference comes only from the fact that the rule for swapping $x$ and $y$ is $(-1)^{|x||y|}$, and we can see that this produces different results in the two cases.
The first step is now to use this, and rewrite the massless $AdS_{3}$ S-matrix in a way which is even more reminiscent of Sine-Gordon. This is easily achieved by using again the proof which we went through in the previous subsection, since we can just reinstate the Sine-Gordon notation while keeping track of the fermionic number:
  • Fendley-Intriligator:
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.13)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{-2H}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{2H}\otimes\hat{F},\quad|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i.$
  • Right-right massless $AdS_{3}\times S^{3}\times T^{4}$:
    $S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (2.14)
    $\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E},\quad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F}$,
    $|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i.$
Rewriting ordinary supersymmetry
The last statement can be seen for instance considering this argument: take for instance
$\Delta\hat{E}=\hat{E}\otimes\mathbbmss{1}+{q_{0}}^{2H-\mathbbmss{1}}\otimes\hat{E}$. (2.15)
This means
$\Delta\hat{E}|\alpha\rangle\otimes|\psi\rangle=(\hat{E}\otimes\mathbbmss{1}+q^{2H-\mathbbmss{1}}\otimes\hat{E})|\phi\rangle\otimes|\psi\rangle=e|\alpha\rangle\otimes|\psi\rangle+(-1)^{\sigma_{3}-\mathbbmss{1}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle,\qquad\alpha=\phi,\psi.$
Both $|\phi\rangle=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $|\psi\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$ have fermionic number $|\phi|=|\psi|=0$, and $\hat{E}$ has zero fermionic number - because it sends the state $|\psi\rangle$ into the state $|\phi\rangle$. We can now rephrase (2.7) in this way:
$\Delta(\hat{E})|\phi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\phi\rangle\otimes|\psi\rangle=x_{2}\,|\phi\rangle\otimes|\phi\rangle+\hat{E}|\phi\rangle\otimes|\psi\rangle,$
$\Delta(\hat{E})|\psi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\psi\rangle\otimes|\psi\rangle=-x_{2}\,|\psi\rangle\otimes|\phi\rangle+\hat{E}|\psi\rangle\otimes|\psi\rangle.$
First, $\hat{E}$ is represented as $x_{i}E_{\phi\psi}=e^{\frac{\theta_{i}}{2}}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$, and has fermionic number $|\hat{E}|=1$; second, $|\phi\rangle$ is represented as $\begin{pmatrix}1\\ 0\end{pmatrix}$, and has fermionic number $|\phi|=0$; third, $|\psi\rangle$ is represented as $\begin{pmatrix}0\\ 1\end{pmatrix}$, and has fermionic number $|\psi|=-1$. The other cases work out analogously. Basically this time the quantum-group deformation $q^{2H-\mathbbmss{1}}$, resp. $q^{-2H+\mathbbmss{1}}$, in the coproduct at this value of the coupling become equal to $(-1)^{|\hat{E}|\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}}$, resp. $(-1)^{|\hat{F}|\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}}$, which simulates ordinary supersymmetry. We have summarised the situation in fig. 1.
Figure 1: Setting the coupling to a special value in Sine-Gordon produces the Fendley-Intriligator model, which has the same symmetry as massless $AdS_{3}$ but different assignment of fermionic numbers of the states.
2.1.1 Going fractional
It is convenient to obtain an $R$-matrix which depends on an unspecified fractional fermionic grading in such a way to interpolate between Sine-Gordon and $AdS_{3}$ when q=iq=-i:
Fractional interpolation:
$S_{a}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\pi\theta}{\xi}}(q^{2}-1)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&\frac{-iqe^{-i\pi\alpha}\big{(}e^{\frac{2\pi\theta}{\xi}}-1\big{)}}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&0\\ 0&\frac{iqe^{i\pi\alpha}\big{(}e^{\frac{2\pi\theta}{\xi}}-1\big{)}}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&\frac{e^{\frac{\pi\theta}{\xi}}(q^{2}-1)}{q^{2}-e^{\frac{2\pi\theta}{\xi}}}&0\\ 0&0&0&1\end{pmatrix}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&S_{R}&ie^{-i\pi\alpha}S_{T}&0\\ 0&-ie^{i\pi\alpha}S_{T}&S_{R}&0\\ 0&0&0&1\end{pmatrix}$,
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+e^{-2H(\log[q]-\log[-i])+\mbox{fracto}}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+e^{2H(\log[q]-\log[-i])-\mbox{fracto}}\otimes\hat{F}$,
$|\phi|=0,\qquad|\psi|=0,\quad|\hat{E}|=0,\quad|\hat{F}|=0,\qquad\mbox{fracto}=\begin{pmatrix}\alpha&0\\ 0&\alpha-1\end{pmatrix}$,
with $\hat{E}=e^{\frac{\pi\theta_{i}}{\xi}}E_{\phi\psi}$, $\hat{F}=e^{\frac{\pi\theta_{i}}{\xi}}E_{\psi\phi}$, and where $S_{R}$, $S_{T}$ are the Sine-Gordon entries (2.2). One can verify that
[…]
and
[…]
We notice that $S_{\alpha}$ solves the Yang-Baxter equation (YBE) for all $\alpha$, and is braiding-unitary $S_{\alpha}(\theta)S_{\alpha}(-\theta)=\mathbbmss{1}\otimes\mathbbmss{1}$ if $\Phi(\theta)\Phi(-\theta)=1$. The parameter $(\alpha,\alpha-1)$ is effectively the statistics of particles $(\phi,\psi)$, which we are using to interpolate between Sine-Gordon (Fendley-Intriligator) $(\frac{1}{2},-\frac{1}{2})$ and $AdS_{3}$ (0,1)(0,-1), as $\alpha$ goes from $\frac{1}{2}$ to 0. We have summarised the situation in fig. 2. $S_{\alpha}$ is also physical-unitary, i.e. $S_{\alpha}$ is a unitary matrix for real rapidities, if $\Phi(\theta)$ is a pure phase for real $\theta$ (and $\alpha$ is real.)
Figure 2: The fractional interpolation connects the diagram 1 which previously was disconnected.
3 Relaxing the special coupling
The basic idea is now to let $q$ be generic in this reformulation of the problem, and see where this leads. The hope is that this introduces the mixed flux in the $AdS_{3}$ case - we mean this purely in terms of functional form of the $S$-matrix, since it is clear that the spectra are different [23] and it is not immediate to us to imagine how they might smoothly deform one into another. So, we could simply recalculate the $S$-matrix with generic $q$, since we can just “forget” that $q$ was special in the above formulas. Applying this to Fendley-Intriligator of course returns back the ordinary Sine-Gordon. For $AdS_{3}$ the result, although potentially interesting per se, is not immediately connected to mixed-flux scattering: this is due to the fact that the mixed-flux relativistic coproduct is twisted in a very particular way [35, 23], so we have to work a bit harder.
There is in fact a way to obtain exactly the mixed flux scattering matrix using this philosophy, by adjusting things in such a way that the deformation of the coproduct is precisely the one observed in the mixed flux case. It works like this. We could have taken the following alternative route to the pure RR massless $AdS_{3}$ $S$-matrix:
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&\epsilon(\theta)&0\\ 0&e^{i\frac{\pi}{k}}\big{(}1-e^{\theta}+e^{i\frac{\pi}{k}+\theta}\epsilon(\theta)\big{)}&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&0\\ 0&0&0&e^{\theta}-e^{i\frac{\pi}{k}}\big{(}1+e^{\theta}\big{)}\epsilon(\theta))\end{pmatrix}$
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}e^{i\pi(q+i)}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}$,
$\qquad|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i,\qquad q=\frac{1}{k}-i,\quad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\quad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
This is obtained using the symmetry alone, and the function $\epsilon(\theta)$ is an arbitrary freedom of the solution not fixed by the symmetry. Solving the Yang-Baxter equation fixes this freedom to
$\epsilon(\theta)=\frac{1}{e^{\frac{i\pi}{k}}+\frac{1}{e^{g}-e^{g+\theta}}},\qquad g\,\,\mbox{constant}.$ (3.1)
We have solved YBE by noticing that it reduces for a single condition for the function $\epsilon(\theta)$ with various arguments. Taking repeated derivatives w.r.t. $\theta_{i}$, $i=1,2,3$, and setting to zero these derivatives, eventually delivers the unique solution (3.1) with a constant of integration remaining arbitrary.
To obtain a limit to the pure RR massless $S$-matrix we can do
$S_{k}\to\lim\limits_{k\to\infty,\epsilon(\theta)\to\tanh\frac{\theta}{2}}\to\,\,\,\,\qquad\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&\tanh\frac{\theta}{2}&0\\ 0&-\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$ (3.2)
$\Delta{\hat{E}}\to\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E},\qquad\Delta{\hat{F}}\to\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F}$.
The choice displayed in the limit (3.2) consistently corresponds to setting $k\to\infty$, and then setting $e^{g}=-\frac{1}{2}$, in (3.1).
We can instead refrain from taking any limit on $q$:
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&\epsilon(\theta)&0\\ 0&e^{i\frac{\pi}{k}}\big{(}1-e^{\theta}+e^{i\frac{\pi}{k}+\theta}\epsilon(\theta)\big{)}&e^{\frac{\theta}{2}}\big{(}1-e^{i\frac{\pi}{k}}\epsilon(\theta)\big{)}&0\\ 0&0&0&e^{\theta}-e^{i\frac{\pi}{k}}\big{(}1+e^{\theta}\big{)}\epsilon(\theta))\end{pmatrix}$
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}e^{i\pi(q+i)}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}$,
$\qquad|\phi|=|\psi|=0,\quad|\hat{E}|=|\hat{F}|=0,\quad q_{0}=-i,\qquad q=\frac{1}{k}-i,\quad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi},\quad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
$\qquad\qquad\to\lim\limits_{\epsilon(\theta)\to-i\frac{\sinh\frac{\theta}{2}}{\sin\big{(}\frac{\pi}{k}-\frac{i\theta}{2}\big{)}}}\to\,\,\,\,\qquad\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&0&0&\frac{\sinh\big{(}\frac{i\pi}{k}-\frac{\theta}{2}\big{)}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}.\end{pmatrix}$ (3.3)
The limit coincides with the left-left relativistic mixed-flux scattering - see for instance [23]. The choice displayed in the limit (3.3) consistently corresponds to setting setting $e^{g}=\frac{i}{2}\mbox{csc}\frac{\pi}{k}$, in (3.1).
To recap, we have an interpolation between pure RR massless and relativistic mixed-flux using the function $\epsilon(\theta)$, which solves the YBE all the way in-between:
$S_{k}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\theta}{2}}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}&\frac{1}{e^{\frac{i\pi}{k}}+\frac{1}{e^{g}-e^{g+\theta}}}&0\\ 0&\frac{e^{i\frac{\pi}{k}}\big{(}1-e^{\theta}\big{)}\big{(}1+e^{i\frac{\pi}{k}+g}\big{)}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}&\frac{e^{\frac{\theta}{2}}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}&0\\ 0&0&0&\frac{1+e^{\theta}}{1-e^{i\frac{\pi}{k}+g}\big{(}e^{\theta}-1\big{)}}-1\end{pmatrix}$ (3.4)
pure RRinterpolationmixed flux
[…]
$k\to\infty,\,\,e^{g}=-\frac{1}{2},\qquad\qquad\qquad\qquad\qquad k\,\,\mbox{generic},\,\,\,e^{g}=\frac{i}{2}\mbox{csc}\frac{\pi}{k}.$ (3.5)
We have summarised the situation in fig. 3. Notice also that $S_{k}$ satisfies braiding unitarity provided $\Phi(\theta)\Phi(-\theta)=1$. It is also physical-unitary if $\Phi(\theta)$ is a pure phase for real $\theta$ and real $k$, and if a certain condition holds on $g$, specifically
$\mbox{Im}(g)=-i\log\Big{[}-\frac{1}{2}e^{-\mbox{Re}(g)-\frac{i\pi}{k}}\Big{(}1+\sqrt{1-4e^{2\mbox{Re}(g)}}\Big{)}\Big{]},$ (3.6)
which is real as long as $\mbox{Re}(g)\geq-\log 2$. The latter condition is satisfied (at least) for both the pure RR and the mixed-flux values of $g$.
Figure 3: $S_{k}$ interpolates between pure RR massless $AdS_{3}$ and relativistic mixed flux.
In mathematical terms, we have basically introduced a Drinfeld twist, which can be formalised for example as:
$R_{k}\to\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}e^{i\pi(q+i)}\otimes\hat{E}=T(\hat{E}\otimes\mathbbmss{1}+q_{0}^{2H-\mathbbmss{1}}\otimes\hat{E})T^{-1}\leftarrow T\bar{R}T^{-1},$
$\qquad\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}e^{-i\pi(q+i)}\otimes\hat{F}=T(\hat{F}\otimes\mathbbmss{1}+q_{0}^{-2H+\mathbbmss{1}}\otimes\hat{F})T^{-1}$
$\qquad T=e^{i\frac{\pi(q+i)}{2}\mathbbmss{1}\otimes\sigma_{3}},\qquad q=\frac{1}{k}-i,$ (3.7)
where $\sigma_{3}=E_{11}-E_{22}$. From the general theory of Drinfeld twists one obtains the relation
$R_{k}=PS_{k},\qquad P=E_{ij}\otimes E_{ji},\qquad T_{21}^{-1}R_{k}T=\bar{R}.$ (3.8)
Similar considerations are at the basis of [54], see also [4] and section 4.3 of [32] for example. Notice that one can use either the $S$-matrix or the associated $R$-matrix - we actually use the appropriate $R$-matrices throughout the paper to prove the Yang-Baxter equation.
3.1 Fractional route
It is again interest to obtain an $S$-matrix which depends on a (more general) unspecified fractional fermionic in such a way to interpolate between Sine-Gordon and $AdS_{3}$ when q=iq=-i, but then to obtain the mixed flux $S$-matrix in another regime.
More general fractional interpolation:
$S_{a+}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{i\pi(\alpha+\beta)}{2}}\big{(}-1+e^{\frac{\theta}{2}}\kappa(\theta)\big{)}}{\sigma}&\kappa(\theta)&0\\ 0&\kappa(\theta)&\sigma e^{\frac{-i\pi(\alpha+\beta)}{2}}\big{(}-1+e^{-\frac{\theta}{2}}\kappa[\theta]\big{)}&0\\ 0&0&0&-1+2\kappa(\theta)\cosh\frac{\theta}{2}\end{pmatrix}$
subject to
$\sigma^{2}=1,\qquad$ (3.9)
$\Delta{\hat{E}}=\hat{E}\otimes\mathbbmss{1}+e^{-2H(\log[q]-\log[-i])+\mbox{Fracto}}\otimes\hat{E},\qquad\Delta{\hat{F}}=\hat{F}\otimes\mathbbmss{1}+e^{2H(\log[q]-\log[-i])-\mbox{Fracto}}\otimes\hat{F}$,
$|\phi|=0,\qquad|\psi|=0,\quad|\hat{E}|=0,\quad|\hat{F}|=0,\qquad\mbox{Fracto}=\begin{pmatrix}\alpha&0\\ 0&\beta-1\end{pmatrix},\quad,\quad\hat{E}=e^{\frac{\theta_{i}}{2}}E_{\phi\psi}\quad\hat{F}=e^{\frac{\theta_{i}}{2}}E_{\psi\phi},$
and subject to the condition
$e^{i\pi\alpha}=-q^{2}\,e^{i\pi\beta},\qquad q=i\sigma e^{\frac{i\pi}{2}(\alpha-\beta)}.$ (3.10)
The function $\kappa(\theta)$ is again an arbitrary freedom of the symmetry solution, restricted by YBE as we will show shortly.
One can verify that
[…]
and
$S_{a+}\to\mbox{massless $AdS_{3}$ \ \ \ \ \ \ \ if}\qquad q\to-i,\,\,\alpha,\beta\to 0,\qquad\kappa(\theta)=\mbox{sech}\frac{\theta}{2},\qquad\sigma=-1.$ (3.12)
Finally, by choosing
$\sigma=-e^{\frac{i\pi(\alpha+\beta)}{2}-\frac{i\pi}{k}},\qquad\kappa(\theta)=\frac{e^{\frac{\theta}{2}}\big{(}e^{\frac{2i\pi}{k}}\ -1\big{)}}{e^{\frac{2i\pi}{k}+\theta}-1},$ (3.13)
we see that
$S_{a+}\to\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&\frac{\sinh\frac{\theta}{2}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&\frac{i\sin\frac{\pi}{k}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}&0\\ 0&0&0&\frac{\sinh\big{(}\frac{i\pi}{k}-\frac{\theta}{2}\big{)}}{\sinh\big{(}\frac{i\pi}{k}+\frac{\theta}{2}\big{)}}\end{pmatrix}$, (3.14)
exactly the mixed flux result. Given that $\sigma$ is a sign, we must have
[…]
In this case it is still possible to have $\beta=\alpha$ provided that
$\alpha=\beta=(\frac{1}{k}+2n+1,\frac{1}{k}+2n),\qquad\qquad n\in\mathbbmss${Z}, (3.15)
which all the way confirms that there is a fractional picture where the fractional supersymmetry is either
$|\phi|=\frac{1}{k}+1,\qquad|\psi|=\frac{1}{k},$ (3.16)
or
$|\phi|=\frac{1}{k},\qquad|\psi|=\frac{1}{k}-1.$ (3.17)
The general solution interpolates between Fendley-Intriligator, pure RR and mixed flux $AdS_{3}$. We impose once again the Yang-Baxter equation to fix the residual freedom in general fashion. By similar techniques as we have used in the previous section we find that the unique solution is
$\kappa(\theta)=\frac{e^{\frac{\theta}{2}}}{1+2\,d\,e^{\frac{\theta}{2}}\sinh\frac{\theta}{2}},\qquad d\,\ \mbox{constant}.$ (3.18)
The interpolation is therefore
$S_{a+}=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\frac{e^{\frac{\theta}{2}}}{1+d(e^{\theta}-1)}&\frac{(1-d)e^{\frac{i\pi(\alpha+\beta)}{2}}\big{(}e^{\theta}-1\big{)}}{\sigma\big{(}1+d(e^{\theta}-1)\big{)}}&0\\ 0&\frac{de^{\frac{-i\pi(\alpha+\beta)}{2}}\big{(}e^{\theta}-1\big{)}\sigma}{1+d(e^{\theta}-1)}&\frac{e^{\frac{\theta}{2}}}{1+d(e^{\theta}-1)}&0\\ 0&0&0&\frac{d(1-e^{\theta})+e^{\theta}}{1+d(e^{\theta}-1)}\end{pmatrix}$ (3.19)
and it reproduces
pure RR / Fendley-Intriligatorinterpolationmixed flux
[…]
$d\to\frac{1}{2},\,\,(\alpha,\beta)=(\frac{1}{2},\frac{1}{2}),\,\,\sigma=-1,\,\,\mbox{for F.-I. or}\qquad\qquad\sigma=-e^{\frac{i\pi(\alpha+\beta)}{2}-\frac{i\pi}{k}},\,\,d\to\frac{1}{2}\big{(}1-i\mbox{cot}\frac{\pi}{k}\big{)}.$
$d\to\frac{1}{2},\,\,(\alpha,\beta)=(0,0),\,\,\sigma=-1,\,\mbox{for \mbox{$AdS_{3}$}}$ (3.20)
We have summarised the situation in fig. 4. Notice that also $S_{a+}$ satisfies braiding unitarity provided $\Phi(\theta)\Phi(-\theta)=1$. It is also physical-unitary if $\Phi(\theta)$ is a pure phase for real $\theta$, and if a certain condition holds on $d$, specifically
$\mbox{Re}(d)=\frac{1}{2},$ (3.21)
 [^7] Unfortunately no million dollars prize is awarded for this.
(satisfied at least by all the ends of the diagram 3.20) and if $\alpha$ and $\beta$ are real.
Figure 4: $S_{a+}$ interpolates between pure RR massless $AdS_{3}$, Fendley-Intrilgator, and relativistic mixed flux.
Finally, it is possible to make a further deformation by replacing everywhere in this section $\theta_{i}\to\frac{2\pi}{\xi}\theta_{i}$, as all the relations continue to hold.
Acknowledgments
We thank Davide Polvara and Bogdan Stefański, jr., for very interesting discussions. We thank Bogdan Stefański, jr., for the wisdom and guidance through the Landau-Ginzburg land. We thank the organisers of IGST 2025 at King’s for a wonderful conference.
No additional research data beyond the data presented and cited in this work are needed to validate the research findings in this work.
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC-BY) licence to any author-accepted manuscript version arising.